OpenStudy (rainbow_dashie):

What is the fraction 24/88 expressed in lowest terms?

OpenStudy (anonymous):

24/88 = 3/11

OpenStudy (rainbow_dashie):

Which fraction expresses – 30/54 in lowest terms?

OpenStudy (rainbow_dashie):

-30/54

OpenStudy (anonymous):

-5/9 i believe

OpenStudy (rainbow_dashie):

@Joel_the_boss @iambatman @iGreen

OpenStudy (anonymous):

Yeah @marie123

OpenStudy (rainbow_dashie):

Which fraction is equivalent to 5/12 ?

OpenStudy (anonymous):

10/24

OpenStudy (rainbow_dashie):

nope

OpenStudy (anonymous):

20/48?

OpenStudy (rainbow_dashie):

hold on

OpenStudy (rainbow_dashie):

A. 2/3 B. 10/18 C. 20/54 D. 15/36

OpenStudy (anonymous):

15/36

OpenStudy (rainbow_dashie):

Which fractions are equivalent to -15/20 ? Choose all answers that are correct. A. Equations below B. C. D.

OpenStudy (anonymous):

is -3/4 on there?

OpenStudy (rainbow_dashie):

\[a-\frac{ 2 }{ 3 }b \frac{ { 9 } }{ -12 }c\frac{ -12 }{ 16 }d\frac{ -25 }{ 30 }\]

OpenStudy (rainbow_dashie):

@I_Need_Help_With_Home

OpenStudy (anonymous):

It's -12/16 I think, Im not really sure on this one